answersLogoWhite

0

One hundred percent it you toss the coin eight times.

User Avatar

Wiki User

∙ 17y ago

What else can I help you with?

Related Questions

What is the probability of you getting three heads straight for tossing a fair coin three times?

It is 1/8 = 0.125


What is the probability of tossing a coin 3 times and getting 3 heads?

There are 23 or 8 possibilities; one is HHH. So, probability of HHH is 1/8 or 0.125.


What is the probability of spinning a number greater than 5 on a spinner numbered 1-8 and tossing on a coin?

3/8. And the coin tossing is totally irrelevant.


What is the probability of getting three heads in tossing a coin three times?

It is (1/2)3 = 1/8 or 0.125


What is the probability of getting exactly 2 heads of tossing 4 coin?

It is 3/8.


Identify the following probability as theoretical or empirical. After tossing the same coin 10 times you are surprised to find that tails has come up 8 times. You therefore conclude that this coin is?

Hhgh


After tossing the same coin 10 times you are surprised to find that tails has come up 8 times You therefore conclude that this coin is not fair and that the probability of getting tails with this co?

Empirical apex students xp


What is the probability of getting heads three times in a row when tossing a coin?

Ideally, the probability of getting any specific combination of length n is 0.5n = 1/2n. For n = 3, this is 0.125 = 1/8.


What is the probability of landing on an even number when tossing an 8 sided die and tossing a coin and getting heads?

1/2 * 1/2 = 1/4 1/2= probability of landing an even number 1/2 = probability of landing a heads


What is the probability of getting one head when flipping a coin three times?

The probability is 3/8.The probability is 3/8.The probability is 3/8.The probability is 3/8.


How many possible outcomes of tossing a coin 4 times?

16


If you toss a fair coin 10 times and it lands on heads 8 times what is the probability that it will land on heads?

Since it is a fair coin, the probability is 0.5

Trending Questions
Would you expect that the inclusion of a parity bit with each character would change the probability of receiving a correct message? What year segragetion ended? What is the meanings of non critical region in statistics? What does the critical value represent? If a coin is tossed 3 times what is the probability of 3 heads? What was the survival rate for the middle passage? What is the type of data that involves numbers? Which is better a 005 level of significance or 001 level of significance? What is the best thing about being a statistician? Why don't you use a circle graph to show data about how body mass changes height? What is the average head size for a 14 year old? What is the probability of randomly choosing a silver tack or a white tack from a box that contains only 15 blue tacks 5 red tacks 20 white tacks and 8 silver tacks? Do instruments and dials provide raw data in it's original form? Difference between Karl Pearson and spearmans rank order coefficient? A manager of a softball team needs to prepare a batting lineup using her 9 starting players How many different lineups can she make? How many pounds are in 150000 kilograms? What is 8.275 to the nearest hundreth of an ounce? What is the approximate percentage score of less than 140 using the 68-95-99.7 rule if a set of test scores is normally distributed with a mean of 100 and a standard deviation of 20? How many outcomes are possible if a coin is tossed 15 times? What is the bit rate a sample rate of a CD?